OpenStudy (anonymous):

The sum of two numbers, a and b, is 12. Their difference is 8. Find the larger number.

OpenStudy (anonymous):

Mathematically, assuming a is the larger number: \[a+b=12\] \[a-b=8\]

OpenStudy (anonymous):

a + b = 12---> a = 12 - b a - b = 8 a - b = 8 12 - b - b = 8 12 - 2b = 8 12 - 8 = 2b 4 = 2b 2 = b a + b = 12 a + 2 = 12 a = 12 - 2 a = 10 check... a + b = 12 10 + 2 = 12 12 = 12 (correct) or ... a - b = 8 10 - 2 = 8 8 = 8 (correct) ANSWER : Larger number is 10. Smaller number is 2